OpenStudy (natasha18):

If G(x) = 3x + 1, then G -1(1) is 4 1/3 0

OpenStudy (natasha18):

i think it's 1/3

OpenStudy (welshfella):

first find the inverse of G(x)

OpenStudy (ingridflowers):

wrong

OpenStudy (natasha18):

x-1/3

OpenStudy (welshfella):

y = 3x + 1 3x = y - 1 x = (y- 1) / 3 g-1(x) = (x - 1) / 3 now plug in x =1 into the above and evaluate

OpenStudy (welshfella):

you can also write g-1(x) as (1/3)x - 1/3

OpenStudy (natasha18):

so 0

OpenStudy (welshfella):

so g-1(1) = (1/3)*1 - 1/3

OpenStudy (welshfella):

Yes 0 is correct

OpenStudy (welshfella):

To find an inverse find x in terms of th function (y) the replace the y by x in the final step

OpenStudy (welshfella):

* then

OpenStudy (welshfella):

OK?

OpenStudy (natasha18):

ok thanks!

OpenStudy (welshfella):

yw